OpenStudy (anonymous):

which point is on the graph of the equation 3x+6y=27 ? a. (0,9) b. (5,2) c. (1,-5) d. (-5,7)

OpenStudy (anonymous):

is it b ?

OpenStudy (anonymous):

B, Plug it in

OpenStudy (anonymous):

3(5)+6(2)=27 15+12=27 27=27

OpenStudy (anonymous):

Yes that should be right

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

No problem! Don't be afraid to ask me!

jimthompson5910 (jim_thompson5910):

The point (-5,7) is also on 3x+6y=27. Here's why: 3x+6y=27 3(-5)+6(7)=27 ... plug in x = -5 and y = 7, from (-5,7) -15 + 42 = 27 27 = 27